OpenStudy (anonymous):

Find the equation of the line that passes through the pair of points. Write your answer in standard form. (2,6) (-5,-1)

OpenStudy (anonymous):

teresa you should try (x1,y1)=(2,6) (x2,y2)=(-5,-1) gradient=m=(y2-y1)/(x2-x1) then insert value m and (x1,y1) or (x2,y2) into y=mx+c to get the value of c and then. rewrite it by inserting m and c value into y=mx+c =) good luck

OpenStudy (anonymous):

how do i get that into standard form?

OpenStudy (anonymous):

standard form is just like the answer i give at the previous question =)